What is the molar concentration of active disinfectant if she dissolve 5 tablespoon of nacl for the sodium hypochlorite in 3.8 m of water

1
Expert's answer
2021-11-19T08:38:02-0500

1 tablespoon = 15g

= 15×5= 75g

Molar mass of NaCl= 58.44

75/58.44= 1.28mol

Molar mass of H2O= 18.02

3.8×18.02=68.48ml

1.28/0.06848= 18.692